OpenStudy (anonymous):

Bridget has a kitchen canister, in the shape of a cylinder, which holds flour. If the canister is 24 centimeters wide and 26 centimeters tall, what is its approximate surface area in terms of pi?

OpenStudy (anonymous):

2 pi r * h is surface area

OpenStudy (anonymous):

2*pi*12*26 = 624*Pi is surface area of cylender if you wanna add the lid to the area then 624*Pi+pir^2=624*Pi+pi*12^2 624*Pi+144*Pi = 768*Pi